### v1

# A Short Proof of the Poincaré Conjecture

Abstract.
A short, fairly self-contained proof is given of the Poincaré Conjecture.
Key words and phrases:
3-manifolds
2010 Mathematics Subject Classification:
20F65 ( 20E08)
1. Introduction
In 2002 I attempted a proof of the Poincaré Conjecture and put it on my home page. A number of errors were pointed out. At the time I was unable to resolve all of them and came to the conclusion that the approach could not work. Recently I wrote an account of my research [2], particularly relating to Stallings’ Theorem and the accessibility of finitely generated groups and that made me think again about my aborted proof. It now seems to me that the approach was a good one and I have come up with a proof that I think resolves the earlier problems. I think the approach could provide the proof of the Sphere Theorem that I was looking for on the last page of [1].
The proof, then and now, was inspired by the beautiful algorithm of Hyam Rubinstein [4] for recognising the 3-sphere and the proof of this by Abigail Thompson [5].
Perelman gave a proof of the Poincaré Conjecture in 2002.
My understanding of simple closed curves on tetrahedrons has benefitted from correspondence with Sam Shepherd.
2. Patterns in a Tetrahedron
Recall from [1] the definition of a pattern. Let $K$ be a finite $2$-complex with polyhedron $|K|$. A pattern is a subset $P$ of $|K|$ satisfying the following conditions:-
  • (i)
    For each $2$-simplex $\sigma$ of $K$, $P\cap|\sigma|$ is a union of finitely many disjoint straight lines joining distinct faces of $\sigma$.
  • (ii)
    For each $1$-simplex $\gamma$ of $K$, $P\cap|\gamma|$ consists of finitely many points in the interior of $|\gamma|$.
A track is a connected pattern. If two patterns $P$ and $Q$ intersect each $1$-simplex in the same number of points then the patterns are said to be equivalent. Two equivalent disjoint tracks in the same $2$-complex are said to be parallel. We investigate tracks and patterns in a tetrahedron $T$, which we regard as the $2$-skeleton $|\rho^{2}|$ of a $3$-simplex $\rho$. We call a track in $T$ an $n$-track if it intersects $n$ edges.
[…]
Figure 1.
If a pattern is as in Figure 1 then the tracks are all $3$-tracks or $4$-tracks. A pattern in a $3$-manifold is called a normal pattern if the intersection with the boundary of every $3$-simplex is like this.
An $8$-track is shown in Figure 2. The only tracks one can have in a tetrahedron are $n$-tracks where $n=3$ or $n=4m$ for $m=1,2,\dots$.
[…]
Figure 2.
[…]
Figure 3.
Figure 3 shows a $12$-track. A pattern in $T$ can only have two types of track. There can be $3$-tracks, each of which separates one of the corner vertices from the other three vertices. There can be one other parallel set of tracks each of which is a $4n$-track for the same positive integer $n$. Each such track separates the four vertices into two pairs. If $n>0$ is even, it separates $u,v$ from $w,z$ and if $n>1$ is odd it separates $u,z$ from $v,w$.
A track in $T$ is a simple closed curve, which will bound a disc in $|\rho|$. If $M$ is a $3$-manifold and $M$ is triangulated so that $M=|K|$ where $K$ is a finite $3$-complex, then a pattern $P$ in $|K|$ determines a patterned surface $S$ such that for each $3$-simplex $\rho$, $S\cap|\rho|$ consists of disjoint properly embedded discs and $S\cap|K^{2}|=P$. A patterned surface is determined, up to isotopy, by the intersection $P\cap|K^{1}|$. If the pattern in $|K^{2}|$ is normal, then the patterned surface is a normal surface.
Orient a track in $T$ by choosing a positive direction as one goes along the track. At adjacent points of intersection of a track with an edge the directions of the track will be opposite to each other. This gives what we call a + - pair, i.e. a pair of points - not necessarily adjacent - of intersection points of an edge with a track where the track has opposite directions. This will be of more significance for singular “tracks ” or stracks as shown in Figures 4 and 5, as a pair of points of intersection on an edge can be removed by a homotopy if and only if the pair is a + - pair. A track in $T$ has a + - pair if and only if it is not a $3$-track or a $4$-track, since any track which intersects an edge more than once will have a + - pair.
A spattern $sP$ in $K$ is defined to be a subset of $|K|$ satisfying
  • (i)
    For each $2$-simplex $\sigma$ of $K$, $sP\cap|\sigma|$ is a union of finitely many straight lines joining distinct faces of $\sigma$.
  • (ii)
    For each $1$-simplex $\gamma$ of $K$, $sP\cap|\gamma|$ consists of finitely many points in the interior of $|\gamma|$. Each such point belongs to exactly one straight line in each of the $2$-simplexes containing $\gamma$.
A strack is a spattern that has no proper subspatterns. Every spattern is a union of finitely many stracks. A strack in $T$ is the image of a circle. If $M$ is a $3$-manifold and $M$ is triangulated so that $M=|K|$ where $K$ is a $3$-complex, then a spattern $sP$ in $|K^{2}|$ determines a spatterned surface $S$ such that for each $3$-simplex $\rho$, $S\cap|\rho|$ consists of singular discs and $S\cap|K^{2}|=sP$.
Let $f:S^{2}\rightarrow M$ be a general position map (see Hempel [3], Chapter 1), in which $f$ is in general position with respect to a triangulation $K$ of $M$. An $i$-piece of $f$ is defined to be a component of $f^{-1}(\sigma)$ where $\sigma$ is an $(i+1)$-simplex of $K$. Thus a 00-piece is a point of $S^{2}$. A $1$-piece is either an scc (simple closed curve) or an arc joining two 00-pieces. If there are no $1$-pieces that are scc’s, then each $2$-piece is a disc whose boundary is a union of $1$-pieces. One can use surgery along simple closed curves to change $f$ to a map in which there are no $1$-pieces that are scc’s, and in which every $2$-piece is a disc. The $2$-pieces will then give a cell decomposition (tessellation) of the $2$-sphere.
If $P$ is a $1$-piece with end points $u,v$ whose images under $f$ are in the same $1$-simplex, then the restriction of $f$ to $P$ is called a returning arc.
In [1] there is a mistake on page 253 of that section on simplifying surface maps. It is incorrectly asserted there that any pair of points in the intersection of the boundary of a $2$-piece with a $1$-simplex can be removed by a homotopy. In fact the pair needs to be a + - pair. Suppose $P$ is $2$-piece for which there is a + - pair $u,v$ in the strack $\delta P$ and that $\gamma$ is the $1$-simplex containing this pair. Let $s:S^{1}\rightarrow\delta P$ be as in the definition of a strack. Possibly by choosing a different + - pair in $\gamma$ we can assume that there is an arc $I=[p,q]\subset S^{1}$ such that $s(p)=u$ and $s(q)=v$ and $s(I)$ intersects $\gamma$ only in its end points $u,v$. There will be a map $s^{\prime}:[p,q]\rightarrow P$ which is close to $s$, for which $s^{\prime}(p)=u,s^{\prime}(q)=v$ and the image of the open interval $(p,q)$ is contained in the interior of the $3$-simplex $\rho$ containing $P$. Any two maps of $I$ into the interior of $\rho$ are homotopic. In particular there will be such a map that is close to the interval in $\gamma$ joining $p,q$. By adjusting $f:S^{2}\rightarrow M$ by such a homotopy and then by another homotopy in a neighbourhood of $I$, we can get a new map $f$ in which the pair $u,v$ has been removed. This is illustrated in Figure 4, where it is shown how the removal of a returning arc will create returning arcs in two other simplexes that have the same $1$-face.
[…]
Figure 4.
Also the pair can be removed without disturbing any other points of intersection with the $1$-skeleton.
[…]
Figure 5.
Figure 5 shows a 12-strack in $T$ that has no + - pairs of points. I conjecture that a strack in $T$ must intersect all six edges to have a + - pair.
If $S$ is a spattern in a $2$-complex $K$, then there is a uniquely determined underlying pattern $P$ that has the same intersection with the $1$-skeleton of $K$. Put $W=S\cap|K^{1}|=P\cap|K^{1}|$.
Figure 6 shows a spattern in $T$ that is a union of a red 12-track and a blue 3-track, and its underlying pattern, which is also a $3$-track and a $12$-track.
[…]
Figure 6.
Given two finite subsets $F_{1},F_{2}$ of the closed interval $[0,1]$ and a bijection $\beta_{I}:F_{1}\rightarrow F_{2}$, there is a continuous map $\phi_{I}:I\rightarrow I$ which restricts to $\beta$ on $F_{1}$ and to the identity on $\{0,1\}$. There is a homotopy between $\phi_{I}$ and the identity map.
Building up from such maps, if $\nu:W\rightarrow W$ is a permutation that restricts to a permutation on $W\cap|\gamma|$ for each $1$-simplex $\gamma$, then $\nu$ extends to a map of the $1$-skeleton into itself which restricts to the identity on the 00-skeleton and which is homotopic to the identity map on $K^{(1)}$. This map can be further extended linearly to a map of the $2$-skeleton and then, in the case of a triangulated $3$-manifold, to a continuous map $\nu:M\rightarrow M$ which is homotopic to the identity map on $M$. It will have the property that if two points on the boundary of a $2$-simplex $\sigma$ are joined by a line in $S\cap\sigma$, then they are joined by a line in $\nu S\cap\sigma$. A spattern $S$ in $K$ is mapped to another spattern. Lines that were uncrossed may become crossed, and lines that were crossed may become uncrossed. The underlying pattern $P$ is not changed. The tessellation of $S$ together with the pieces of the tessellation are unchanged in such a map.
Our proof of the Poincaré Conjecture is to show that a certain spattern must occur in a homotopy from the boundary of a fake ball to a constant map, and this spattern is homotopic to its underlying pattern by such a homotopy.
3. The Proof
Let $M=|K|$ be a $3$-manifold, where $K$ is a finite $3$-complex. It follows from a result of Kneser (see [1] or [3]) that there is a bound on the number of disjoint non parallel normal surfaces in a compact triangulated $3$-manifold. I was able to prove that finitely presented groups are accessible by generalising this result to patterns in a finite $2$-complex. In the Recognition Algorithm one determines a maximal set of disjoint normal surfaces in a triangulated 3-manifold M that are 2-spheres. If $M$ is simply connected, then each such surface separates M and so the set of surfaces correspond to the edges of a finite tree. It is proved that M is a 3-sphere if each region corresponding to a vertex of valency one in this tree either contains a single vertex of the triangulation or contains no vertices but does contain an almost normal surface, i.e. one for which the intersections of the surface with 3-simplexes are all 3 or 4-sided except for one exceptional 8-sided disc.
Let $M$ be a fake 3-ball. Let $M_{0}$ be a component, obtained by cutting along the maximal collection of normal $2$-spheres, which has one boundary component and which does not contain a vertex. Let $f:S^{2}\rightarrow M$ be an injective map whose image is the normal $2$-sphere $\delta M_{0}$. Since $M_{0}$ is a fake $3$-ball, the map $f$ is null homotopic. Let $F:S^{2}\times I\rightarrow M_{0}$ be a homotopy between $f$ and a constant map. For $t\in I$ let $f_{t}:S^{2}\rightarrow M_{0},f_{t}(s)=F(s,t)$. We can assume that for all but finitely many values $t,f_{t}$ meets the $1$-skeleton $T^{1}$ of the triangulation transversely and for each $t$ for which the map $f_{t}$ does not meet $T^{1}$ transversely, there is precisely one point where $f_{t}$ is tangential to $T^{1}$. Let $t_{1}^{\prime},t_{2}^{\prime},\dots,t_{n}^{\prime}$ be the values of $t$ for which $f_{t}$ does not meet $T^{1}$ transversely and put $t_{0}^{\prime}=0,t_{n+1}^{\prime}=1$. For $i=0,1,...,n$ choose $t_{i}\in(t_{i}^{\prime},t_{i+1}^{\prime})$ and put $f_{i}=f_{t_{i}}$.
Let $W_{i}$ be the set of intersections of $f_{i}(S^{2})$ with the $1$-skeleton of $T$ and let $w_{i}=|W_{i}|$. Rearrange the weights $w_{i}$ into a finite non-increasing sequence. Order these sequences lexicographically. The width of $T$ is the minimum sequence of weights as $F$ ranges over all possible homotopies. If F realizes the width of $T$, then $F$ is said to be in thin position. Now revert to the original ordering of the $w_{i}$’s. Clearly for each $i$ either $w_{i+1}=w_{i}+2$ or $w_{i+1}=w_{i}-2$. Note that $w_{0}$ is the number of intersections of $\delta M_{0}$ with the $1$-skeleton. For $F$ in thin position, we are particularly interested in the values $i$ for which $w_{i-1}=w_{i+1}=w_{i}-2$. For such an $i,f_{i}(S^{2})$ is called a thick sphere, while if $w_{i-1}=w_{i+1}=w_{i}+2$, then $f_{i}(S^{2})$ is called a thin sphere. Our interest will be in the first thick sphere. Let $S=f_{k}(S^{2})$ be the first thick sphere.. Each of the preceding $f_{j}(S^{2})=S_{j}$, $1\leq j\leq k$, satisfies $w_{j-1}=w_{j}-2$, so that $W_{j}=W_{j-1}\cup\{u_{j},v_{j}\}$ where $u_{j},v_{j}$ are a pair of points, labelled $j,j$, from a particular $1$-simplex $\gamma_{j}$.
(a)(b)
Figure 7.
Now consider $f_{k}:S^{2}\rightarrow M_{0}$ with $S=f_{k}(S^{2})$. We follow the argument of [5]. We know that the homotopy going from $f_{k}$ to $f_{k+1}$ results in the deletion of two points $x,y$ in a $1$-simplex $\gamma$ and the homotopy going from $f_{k}$ to $f_{k-1}$ deletes points $u,v$. We also know that these pairs must be + - pairs. All four points must belong to the same $2$-piece, for if there is no $2$-piece that contains all four points, then the order of the homotopies can be changed and the weight of the total homotopy reduced - one peak is replaced by two smaller peaks. Also if $x,y,u,v$ are in the same $2$-piece and $x,y$ are still in the same $2$-piece for the map $f_{k-1}$, then we can also swap the homotopies round and reduce weight. Similarly if $u,v$ are in the same $2$-piece for the map $f_{k+1}$, then we can swap the homotopies. It follows that the exceptional $2$-piece has two pairs of vertices that can be removed by a homotopy and that removing one pair disconnects the $2$-piece. The simplest such $2$-piece has boundary an $8$-track as in Figure 2 or Figure 7(a). Note that $x,y$ will be two points on one edge and ${u,v}$ will be the vertices on the opposite edge. Removing $u,v$ (labelled $k,k$) creates two 3-gons, with $x,y$ separated as shown in Figure 8. In an isotopy this is the only possibility for the exceptional $2$-piece. For a homotopy there are other possibilities such as in Figure 7(b).
[…]
Figure 8.
We will show that all the $2$-pieces apart from the exceptional one have no + - pairs of vertices which map to the same $1$-simplex. This is similar to the proof in [5]. Suppose $p,q$ are such a pair of vertices in a $2$-piece. These can be removed by a homotopy. Suppose first that $p,q$ are both different from any of $u,v,x,y$. Starting from the $(k-1)$-th stage, we can carry out a homotopy that removes $p,q$. We then add in the pair $u,v$ and remove $x,y$ and finally replace $p,q$ so that we are at the $(k+1)$-th stage. This sequence of homotopies has a lower sequence of weights than the original as the weight of the highest peak has been reduced by two. If, say $p=u$ but $q\not=v$, then we can remove $p,q$. This will not separate $x,y$ and so we can remove $x,y$ and then add in $u,v$. This has changed the points of $W_{k+1}$ by replacing $q$ by $v$ and this can be achieved by a homotopy which only affects the points on one $1$-simplex. Again this sequence of homotopies has a lower weight sequence.
Let $W=W_{k}$. It is clear that $S=f_{k}(S^{2})$ has no returning arcs and so it is a spattern. Let $P$ be the underlying pattern. It will be shown that there is a permutation of $W$, restricting to a permutation on each intersection with a $1$-simplex for which the corresponding homotopy changes $S$ to $P$.
We see that $W_{0}$ is the set of vertices of a normal $2$-sphere $S_{0}$ parallel to the normal $2$-sphere $\delta M_{0}$. The pair $u_{1},v_{1}$ will be the ends of a returning arc in at least one $2$-simplex. In another $2$-simplex $u_{1},v_{1}$ will be joined by lines in both $S_{1}$ and $S$ to the vertices of an edge in $S_{0}$.
$u_{1}$$v_{1}$$x$$y$$w$$z$
Figure 9.
Figure 9 shows the situation for $u_{1},v_{1}$ which lie in the $1$-simplex $\gamma_{1}$. In at least one $2$-simplex containing $\gamma_{1}$ as a face, $u_{1},v_{1}$ will be joined to the vertices $w,z$ of an edge of $S_{0}$, shown dashed red. In $P$, $w,z$ will be joined to points $x,y$ by lines, shown blue in Figure 9. Let $\nu_{1}$ be the permutation of $W$, which restricts to $(v_{1},x)(u_{1},y)$ for $W\cap\gamma_{1}$ and is the identity map on all the other $1$-simplexes. Under the corresponding homotopy the two lines become uncrossed if they were crossed before. Note that the end points are adjacent and stay adjacent because the homotopy takes place in $M_{0}$ and so the new lines of the homotopy do not intersect the lines of $S_{0}$. The pattern for $S_{0}$ in each $3$-simplex is as in Figure 1 and so the homotopy $\nu_{1}$ must take place in one of the regions which is not bounded by two parallel lines. This is illustrated in Figure 10, which indicates what can happen in the faces of a $3$-simplex. If a $2$-simplex $\sigma$ contains $\gamma_{1}$ as one of its faces, but has no intersection with $S_{0}$, then the points will be joined by a returning arc in $\sigma$.
[…]
Figure 10.
We use an induction argument for defining $\nu_{i}$ for $2\leq i\leq k$.
$u_{i}$$v_{i}$$x$$y$$u_{i}$$v_{i}$$x$$y$(a)(b)
Figure 11.
Our aim is to show that for each $1$-simplex $\gamma$ we can permute the finite set $\gamma\cap S$ in such a way that under the associated homotopies the spattterned $2$-sphere $S$ becomes the underlying patterned $2$-sphere $P$. It will then be the case that $F$ becomes an isotopy.
Let $\gamma_{i}$ be the $1$-simplex containing $u_{i}$ and $v_{i}$.
We have defined $\nu_{1}$. Let $\mu_{1}=\nu_{1}$. We now define $\nu_{i}$ and $\mu_{i}$ for $i=2,\dots,k$ to be permutations of $W$, where $\nu_{i}$ restricts to a permutation on the points of $W\cap\gamma_{i}$ and is the identity on the other points of $W$, and so that the homotopy associated with $\mu_{i}=\nu_{i}\mu_{i-1}$ moves the points of $W_{i}$ to the positions they should have in $P$. In at least one of the $2$-simplexes containing $\gamma_{i}$ as a face there are lines in $S$ joining $u_{j}$ to $u_{i}$ for some $j<i$ or there is a situation as in Figure 9, in which $u_{i},v_{i}$ are joined to the vertices of an edge of $S_{0}$.
If there are lines in $\mu_{i-1}S$ joining $u_{i}$ to $u_{j}$ and $v_{i}$ to $v_{j}$ it will be as in Figure 11(a) or there will be three labelled pairs joined by lines as in Figure 11(b). This is because $S_{i}$ has a returning arc joining $u_{i}$ and $v_{i}$ and this returning arc was created by one or more homotopy $f_{j}$ for $j<i$.
In Figure 11 (a) we see what happens in a $2$-simplex $\sigma_{i}$ containing $\gamma_{i}$, if the lines joining $u_{i}$ and $v_{i}$ come from a single pair $u_{j},v_{j}$. In Figure 11(b) the lines to $u_{i},v_{i}$ come from two pairs. The lines in blue are those of $\mu_{i-1}S\cap\sigma_{i}$. The lines in red are the ones we move them to in $\mu_{i}S$. The bottom edge is the $1$-simplex $\gamma_{i}$ containing the pair $u_{i},v_{i}$ which have labels $i,i$ in the labelling for $S$. In at least one of the $2$-simplexes containing $\gamma$ there will be a $1$-face different from $\gamma$ containing a pair of adjacent vertices with labels $l,l$ for $l<i$, which are joined by non-intersecting lines in $P$ to the pair $x,y$ in $\gamma_{i}$, or the situation will be as in Figure 11(b) so that there are pairs with labels less than $i$ in each of the other two faces. The induction step is to apply the permutation $\nu_{i}=(v_{i},x)(u_{i},y)$ to the set $\gamma_{i}\cap W$. Of course if $u_{i}=y$ we apply the transposition $(v_{i},x)$. Again with each transposition we also transpose all attaching lines in all $2$-simplexes containing $\gamma$. This change will not disturb any of the vertices of $W_{i-1}$ or the edges joining them. Note that $x,y$ will have labels bigger than $i$ as if, say, $x\in W_{i-1}$, then in $S_{i-1}$ there would be a point in the intersection of a $1$-simplex with $\mu_{i-1}S$ joined to both $u_{i}$ and $y$ and this does not happen in a spattern..
It now follows that at the end of the induction $P=\mu_{k}S$ so that $P$ determines a patterned $2$-sphere.
Put $\mu=\mu_{k}$. We now know that $\mu S$ is a patterned surface. All the $2$-pieces, apart from the exceptional one, intersect each $1$-simplex at most once and so are $3$-sided or $4$-sided. The $8$-track shown in Figure 6(a) is the only possibility for the exceptional $2$-piece. Thus $S$ has become an almost normal $2$-sphere and we have a proof of the Poincaré Conjecture. In the homotopy $F$, the first thick sphere $S$ will also be the last thick sphere. After applying $\mu$, then for $j>k$ each step of the homotopy becomes an isotopy in which a returning arc joining adjacent points is removed.