### v3

# A Short Proof of the Poincaré Conjecture

Abstract.
A short, fairly self-contained proof is given of the Poincaré Conjecture. There is a possible error on Page 8. Hopefully it may just be a gap that can be filled.
Key words and phrases:
3-manifolds
2010 Mathematics Subject Classification:
20F65 ( 20E08)
1. Introduction
In 2002 I attempted a proof of the Poincaré Conjecture and put it on my home page. A number of errors were pointed out. At the time I was unable to resolve all of them and came to the conclusion that the approach could not work. Recently I wrote an account of my research [2], particularly relating to Stallings’ Theorem and the accessibility of finitely generated groups and that made me think again about my aborted proof. It now seems to me that the approach was a good one and I have come up with a proof that I think resolves the earlier problems. I think the approach could provide the proof of the Sphere Theorem that I was looking for on the last page of [1].
The proof, then and now, was inspired by the beautiful algorithm of Hyam Rubinstein [4] for recognising the 3-sphere and the proof of this by Abigail Thompson [5]. In fact we show that the argument of [5] can be generalised to apply to a sequence of homotopies rather than a sequence of isotopies.
Perelman gave a proof of the Poincaré Conjecture in 2002.
My understanding of simple closed curves on tetrahedrons has benefitted from correspondence with Sam Shepherd and with Andy Bartholomew.
2. Patterns in a Tetrahedron
Recall from [1] the definition of a pattern. Let $K$ be a finite $2$-complex with polyhedron $|K|$. A pattern is a subset $P$ of $|K|$ satisfying the following conditions:-
  • (i)
    For each $2$-simplex $\sigma$ of $K$, $P\cap|\sigma|$ is a union of finitely many disjoint straight lines joining distinct faces of $\sigma$.
  • (ii)
    For each $1$-simplex $\gamma$ of $K$, $P\cap|\gamma|$ consists of finitely many points in the interior of $|\gamma|$.
A track is a connected pattern. If two patterns $P$ and $Q$ intersect each $1$-simplex in the same number of points then the patterns are said to be equivalent. Two equivalent disjoint tracks in the same $2$-complex are said to be parallel. We investigate tracks and patterns in a tetrahedron $T$, which we regard as the $2$-skeleton $|\rho^{2}|$ of a $3$-simplex $\rho$. We call a track in $T$ an $n$-track if it has $n$ intersections with edges.
[…]
Figure 1.
If a pattern is as in Figure 1 then the tracks are all $3$-tracks or $4$-tracks. A pattern in a $3$-manifold is called a normal pattern if the intersection with the boundary of every $3$-simplex is like this.
An $8$-track is shown in Figure 2. The only tracks one can have in a tetrahedron are $n$-tracks where $n=3$ or $n=4m$ for $m=1,2,\dots$.
[…]
Figure 2.
[…]
Figure 3.
Figure 3 shows a $12$-track. A pattern in $T$ can only have two types of track. There can be $3$-tracks, each of which separates one of the corner vertices from the other three vertices. There can be one other parallel set of tracks each of which is a $4n$-track for the same positive integer $n$, and for which there is a fixed sum $n=a+b$, where $a,b$ are coprime integers. If $n$ is even, then $a,b$ are both odd integers, and the track is as in Figure 2 with $a$ lines going from $uv$ to $wz$ and each of the other $4$ lines replaced by $b$ parallel lines. If $n$ is odd, then exactly one of $a,b$ is even. If $a$ is even, so that $b$ is odd, then the $4n$-track is as in Figure 3 again with $a$ lines going from $uv$ to $wz$ and $b$ parallel lines replacing each of the other $4$ lines. These are patterns because the numbers of intersection points on edges match up. Each is a track because any proper subpattern must be of the same form with smaller integers $a,b$. Each such pattern separates the four vertices and the four centre points into two pairs. If $a$ is odd as in Figure 2, the $4n$-pattern separates $u,v$ from $w,z$ and $e,f$ from $g,h$. While if $a$ is even as in Figure 3, the $4n$-track separates $u,z$ from $v,w$ and $e,g$ from $f,h$. Thus each such subpattern will separate vertices and centre points in the same way and the space between two component tracks of the pattern could contain no vertex or centre point, so that in the terminology of [1] it is an untwisted band, making the two tracks parallel. But $a,b$ will not be coprime if all the tracks in the pattern are parallel and there is more than one. Thus each pattern is a $4n$-track.
In fact for our proof here, we only need to know that there is an $8$-track as in Figure 2 that intersects each of two opposite edges of $T$ in two points and each of the other four edges just once.
A track in $T$ is a simple closed curve, which will bound a disc in $|\rho|$. If $M$ is a $3$-manifold and $M$ is triangulated so that $M=|K|$ where $K$ is a finite $3$-complex, then a pattern $P$ in $|K^{2}|$ determines a patterned surface $S$ such that for each $3$-simplex $\rho$, $S\cap|\rho|$ consists of disjoint properly embedded discs and $S\cap|K^{2}|=P$. A patterned surface is determined, up to isotopy, by the intersection $P\cap|K^{1}|$. If the pattern in $|K^{2}|$ is normal, then the patterned surface is a normal surface.
Orient a track in $T$ by choosing a positive direction as one goes along the track. At adjacent points of intersection of a track with an edge the directions of the track will be opposite to each other. This gives what we call a + - pair, i.e. a pair of points - not necessarily adjacent - of intersection points of an edge with a track where the track has opposite directions. This will be of more significance for singular “tracks ” or stracks as shown in Figures 5 and 6, as if a pair of points of intersection on an edge can be removed by a homotopy, then the pair is a + - pair. A track in $T$ has a + - pair if and only if it is not a $3$-track or a $4$-track, since adjacent points of intersection of a track with an edge will be a + - pair.
A spattern $sP$ in $K$ is defined to be a subset of $|K|$ satisfying
  • (i)
    For each $2$-simplex $\sigma$ of $K$, $sP\cap|\sigma|$ is a union of finitely many straight lines joining distinct faces of $\sigma$.
  • (ii)
    For each $1$-simplex $\gamma$ of $K$, $sP\cap|\gamma|$ consists of finitely many points in the interior of $|\gamma|$. Each such point belongs to exactly one straight line in each of the $2$-simplexes containing $\gamma$.
A strack is a spattern that has no proper subspatterns. Every spattern is a union of finitely many stracks. A strack in $T$ is the image of a circle. If $M$ is a $3$-manifold and $M$ is triangulated so that $M=|K|$ where $K$ is a $3$-complex, then a spattern $sP$ in $|K^{2}|$ determines a spatterned surface $S$ such that for each $3$-simplex $\rho$, $S\cap|\rho|$ consists of singular discs and $S\cap|K^{2}|=sP$.
Let $f:S^{2}\rightarrow M$ be a general position map (see Hempel [3], Chapter 1), in which $f$ is in general position with respect to a triangulation $K$ of $M$. An $i$-piece of $f$ is defined to be a component of $f^{-1}(\sigma)$ where $\sigma$ is an $(i+1)$-simplex of $K$. Thus a 0-piece is a point of $S^{2}$. A $1$-piece is either an scc (simple closed curve) or an arc joining two 0-pieces. If there are no $1$-pieces that are scc’s, then each $2$-piece has boundary that is is a union of $1$-pieces. One can use surgery along simple closed curves to change $f$ to a map in which there are no $1$-pieces that are scc’s, and in which every $2$-piece is a disc. The $2$-pieces will then give a cell decomposition (tessellation) of the $2$-sphere.
If $R$ is a $1$-piece with end points $u,v$ whose images under $f$ are in the same $1$-simplex, then the restriction of $f$ to $R$ is called a returning arc.
If $f:S^{2}\rightarrow M$ has no $1$-pieces that are returning arcs or scc’s, then the intersection of $f(S^{2})$ with the $2$-skeleton of $M$ is a spattern $sP$. In the case in which we are interested, there is a homotopy from $f$ to $f^{\prime}:S^{2}\rightarrow M$ in which the image is the spatterned surface determined by $sP$.
Let $\gamma$ be a $1$-simplex of $M$. Two points $p,q\in\gamma\cap f(S^{2})$ are said to be removable if there is a homotopy from $f$ to a map $f^{\prime}:S^{2}\rightarrow M$ such that $f(x)=f^{\prime}(x)$ for every $x$ that is not in the interior of a simplex with $\gamma$ as a face and $\gamma\cap f^{\prime}(S^{2})$ is the same as $\gamma\cap f(S^{2})$ but with $p,q$ removed.
The pair of end points of a returning arc $R$ are removable by the following homotopy. Let $\sigma$ be the $2$-simplex of $K$ such that $f(R)\subset\sigma$. Let $V$ be a regular neighbourhood of $R$ in $S^{2}$. Let $V^{\circ}$ be the interior of $V$ regarded as a subspace of $V$. Let $\beta V$ be the boundary of $V$ regarded as a subspace of $S^{2}$, so that $\beta V=V-V^{\circ}$. Let $\gamma$ be the $1$-simplex containing the end points of $R$. The regular neighbourhood $V$ is a disc and $\beta V=\delta V$ is a simple closed curve in $S^{2}$. The union of all the $3$-simplexes that contain $\gamma$ is a closed ball $B$ and $f(\beta V)\subset B^{\circ}-\sigma$, which is contractible. Define $f^{\prime}:S^{2}\rightarrow M$ so that $f^{\prime}$ is continuous, $f^{\prime}$ and $f$ are the same when restricted to $S^{2}-V^{\circ}$, and $f^{\prime}(V)\subset B^{\circ}-\sigma$. Note that removing $p$ may create more $1$-pieces that are returning arcs or sccs, but the size of the intersection with the $1$-skeleton goes down by two. In the case in which we are interested intersections, in which sccs can be removed as above, the maps $f$ and $f^{\prime}$ are homotopic. This is illustrated in Figure 4, where it is shown how the removal of the ends of a returning arc will create returning arcs in two other simplexes that have the same $1$-face.
There are also removable pairs of points if a strack intersects a $1$-simplex in a + - pair. In [1] there is a mistake on page 253 in the section on simplifying surface maps. It is incorrectly asserted there that any pair of points in the intersection of the boundary of a $2$-piece with a $1$-simplex can be removed by a homotopy.
Suppose $Q$ is a $2$-piece and that $\gamma$ is a $1$-simplex for which $\delta Q\cap\gamma$ contains at least one + - pair. Let $s:S^{1}\rightarrow\delta Q$ be as in the definition of a strack. There will be at least one + - pair in $\gamma$ for which there is an arc $I=[p,q]\subset S^{1}$ such that $s(p)=u$ and $s(q)=v$ and $s(I)$ intersects $\gamma$ only in its end points $u,v$. Such a pair is removable. Thus if $\gamma$ contains a + - pair, then it contains a removable pair. Let $p,q$ be a removable pair as above. There will be a map $s^{\prime}:[p,q]\rightarrow Q$ which is close to $s$, for which $s^{\prime}(p)=u,s^{\prime}(q)=v$ and the image of the open interval $(p,q)$ is contained in the interior of the $3$-simplex $\rho$ containing $Q$. Any two maps of $I$ into the interior of $\rho$ are homotopic. In particular there will be such a map that is close to the interval $J$ in $\gamma$ joining $p,q$. We can adjust $f:S^{2}\rightarrow M$ by a homotopy so that the piece $Q$ contains the image of this map. Thus $f^{\prime}:S^{2}\rightarrow M$ is the same outside $Q$ and on the boundary of $Q$, but in the interior of $Q$ takes the arc close to $s$ to the arc $J$. Now $f$ and $f^{\prime}$ are homotopic. Another homotopy in a neighbourhood of $I$, will give a new map $f^{\prime\prime}$ in which the pair $p,q$ has been removed. This homotopy is similar to the one for removing a returning arc.
$\bullet\bullet-++\bullet\bullet$-
Figure 4.
In both cases, a removable pair can be removed without disturbing any other points of intersection with the $1$-skeleton.
[…]
Figure 5.
Figure 5 shows an $11$-strack in $T$ that has no + - pairs of points.
If $S$ is a spattern in a $2$-complex $K$, then there is a uniquely determined underlying pattern $P$ that has the same intersection with the $1$-skeleton of $K$. Put $W=S\cap|K^{1}|=P\cap|K^{1}|$.
Figure 6 shows a spattern in $T$ that is a union of a red 12-track and a blue 3-track, and its underlying pattern, which is also a $3$-track and a $12$-track. The underlying pattern for the strack of Figure 5 has two $4$-tracks. and one $3$-track.
[…]
Figure 6.
Given two finite subsets $F_{1},F_{2}$ of the closed interval $I=[0,1]$ and a bijection $\beta_{I}:F_{1}\rightarrow F_{2}$, there is a continuous map $\phi_{I}:I\rightarrow I$ which restricts to $\beta$ on $F_{1}$ and to the identity on $\{0,1\}$. There is a homotopy between $\phi_{I}$ and the identity map.
Building up from such maps, if $\nu:W\rightarrow W$ is a permutation that restricts to a permutation on $W\cap|\gamma|$ for each $1$-simplex $\gamma$, then $\nu$ extends to a map of the $1$-skeleton into itself which restricts to the identity on the 0-skeleton and which is homotopic to the identity map on $K^{(1)}$. This map can be further extended linearly to a map of the $2$-skeleton and then, in the case of a triangulated $3$-manifold, to a continuous map $\nu:M\rightarrow M$ which is homotopic to the identity map on $M$. It will have the property that if two points on the boundary of a $2$-simplex $\sigma$ are joined by a line in $S\cap\sigma$, then they are joined by a line in $\nu S\cap\sigma$. A spattern $S$ in $K$ is mapped to another spattern. Lines that were uncrossed may become crossed, and lines that were crossed may become uncrossed. The underlying pattern $P$ is not changed. The tessellation of $S$ together with the pieces of the tessellation are unchanged in such a map.
Our proof of the Poincaré Conjecture is to show that a certain spattern must occur in a homotopy from the boundary of a fake ball to a constant map, and this spattern is homotopic to its underlying pattern by such a homotopy.
3. The Proof
Let $M=|K|$ be a $3$-manifold, where $K$ is a finite $3$-complex. It follows from a result of Kneser (see [1] or [3]) that there is a bound on the number of disjoint non parallel normal surfaces in a compact triangulated $3$-manifold. I was able to prove that finitely presented groups are accessible by generalising this result to patterns in a finite $2$-complex. In the Recognition Algorithm one determines a maximal set of disjoint normal surfaces in a triangulated 3-manifold M that are 2-spheres. If $M$ is simply connected, then each such surface separates M and so the set of surfaces correspond to the edges of a finite tree. It is proved that M is a 3-sphere if each region corresponding to a vertex of valency one in this tree either contains a single vertex of the triangulation or contains no vertices but does contain an almost normal surface, i.e. one for which the intersections of the surface with 3-simplexes are all 3 or 4-sided except for one exceptional 8-sided disc.
Let $M$ be a fake 3-ball. Let $M_{0}$ be a component, obtained by cutting along the maximal collection of normal $2$-spheres, which has one boundary component and which does not contain a vertex. By Van Kampen, $M_{0}$ is simply connected, and so it is either a ball or a fake ball. In either case there is a homotopy between the boundary and the constant map. Let $f:S^{2}\rightarrow M$ be an injective map whose image is the normal $2$-sphere $\delta M_{0}$. Let $F:S^{2}\times I\rightarrow M_{0}$ be a homotopy between $f$ and a constant map. For $t\in I$ let $f_{t}:S^{2}\rightarrow M_{0},f_{t}(s)=F(s,t)$. We can assume that for all but finitely many values $t,f_{t}$ meets the $1$-skeleton $T^{1}$ of the triangulation transversely and for each $t$ for which the map $f_{t}$ does not meet $T^{1}$ transversely, there is precisely one point where $f_{t}$ is tangential to $T^{1}$. Let $t_{1}^{\prime},t_{2}^{\prime},\dots,t_{n}^{\prime}$ be the values of $t$ for which $f_{t}$ does not meet $T^{1}$ transversely and put $t_{0}=0,t_{n+1}=1$. For i=1,,ni=1,...,n choose $t_{i}\in(t_{i}^{\prime},t_{i+1}^{\prime})$ and put $f_{i}=f_{t_{i}}$.
Let $W_{i}$ be the set of intersections of $f_{i}(S^{2})$ with the $1$-skeleton of $T$ and let $w_{i}=|W_{i}|$. Rearrange the weights $w_{i}$ into a finite non-increasing sequence. Order these sequences lexicographically. The width of $T$ is the minimum sequence of weights as $F$ ranges over all possible homotopies. If F realises the width of $T$, then $F$ is said to be in thin position. Now revert to the original ordering of the $w_{i}$’s. Clearly for each $i$ either $w_{i+1}=w_{i}+2$ or $w_{i+1}=w_{i}-2$. Note that $w_{0}$ is the number of intersections of $\delta M_{0}$ with the $1$-skeleton. For $F$ in thin position, we are particularly interested in the values $i$ for which $w_{i-1}=w_{i+1}=w_{i}-2$. For such an $i,f_{i}(S^{2})$ is called a thick sphere, while if $w_{i-1}=w_{i+1}=w_{i}+2$, then $f_{i}(S^{2})$ is called a thin sphere. Our interest will be in the first thick sphere. Since there are no removable pairs in a normal surface, $w_{1}=w_{0}+2$. Let $S=f_{k}(S^{2})$ be the first thick sphere. Each of the preceding $f_{j}(S^{2})=S_{j}$, $1\leq j\leq k$, satisfies $w_{j-1}=w_{j}-2$, so that $W_{j}=W_{j-1}\cup\{u_{j},v_{j}\}$ where $u_{j},v_{j}$ are a pair of points, labelled $j,j$, from a particular $1$-simplex $\gamma_{j}$.
Now consider $f_{k}:S^{2}\rightarrow M_{0}$ with $S=f_{k}(S^{2})$. We follow the argument of [5]. We know that the homotopy going from $f_{k}$ to $f_{k+1}$ results in the deletion a removable pair in a $1$-simplex $\gamma$ and the homotopy going from $f_{k}$ to $f_{k-1}$ results in the removal of another pair. Both pairs must belong to the same $2$-piece, for if there is no $2$-piece that contains both pairs, so that one pair is in one $2$-piece and the other removable pair is in another $2$-piece, then the order of the homotopies can be changed and the weight sequence of the total homotopy reduced - one peak is replaced by two smaller peaks. Also if both pairs are in the same $2$-piece and one pair is not separated by the removal of the other pair, then we can also swap the homotopies round and get a lower weight sequence. It follows that the exceptional $2$-piece has two removable pairs and that removing one pair disconnects the $2$-piece. The simplest such $2$-piece has boundary an $8$-track as in Figure 2. Removing one pair (labelled $k,k$) creates two 3-tracks, with the pair on the opposite edge separated as shown in Figure 7. In an isotopy this is the only possibility for the exceptional $2$-piece. For a homotopy there are other possibilities such as the $12$-strack in Figure 8. In this case removing the pair labelled $k,k$ gives a blue $3$-track and a red $7$-strack. In both cases, note that the lines joining the pair labelled $k,k$ connect to points on the other two edges of the 22–simplex. They are neither returning arcs, nor do they connect to points on the same $1$-simplex. If they did, then removal of the pair labelled $k,k$ would not disconnect the strack.
[…]
Figure 7.
[…]
Figure 8.
We will show that all the $2$-pieces apart from the exceptional one have no removable pairs. This is similar to the proof in [5]. Suppose $p,q$ are a removable pair of 0-pieces in a $2$-piece different from the exceptional $2$-piece. Starting from the (k1)(k-1)-th stage, we can carry out a homotopy that removes $p,q$. We then carry out the homotopies $f_{k}$ and $f_{k+1}$ before replacing $p,q$. This sequence of homotopies has a lower sequence of weights than the original, as the weight of the highest peak has been reduced by two. If $p,q$ are in the exceptional $2$-piece but are also in $W_{k-1}$ or $W_{k+1}$, then we can also construct a sequence of homotopies with a lower weight sequence.
It is clear then that $S=f_{k}(S^{2})$ has no returning arcs. Also it can be assumed that it has no $1$-pieces that are scc’s. For if there was a $1$-piece that was an scc then surgery along this curve would produce two $2$-spheres, one of which will contain the exceptional $2$-piece. If this $2$-sphere $S^{\prime}$ had fewer intersections with the $1$-skeleton, i.e. $S^{\prime}\cap K^{1}$ is a proper subset of $W=W_{k}$, then it would contradict the thin position of $f$. Thus $S^{\prime}\cap K^{1}=W$ and the other $2$-sphere will not intersect the $1$-skeleton. All the $1$-pieces for this $2$-sphere must be scc’s, and starting from an innermost one, these can be removed from $f$ by a homotopy. Using a similar argument one can do surgery along scc’s to change every $2$-piece in $S$ that is not a disc to one that is a disc.
We can assume, then, that $S$ is a spattern. Let $P$ be the underlying pattern. It will be shown that there is a permutation of $W$, restricting to a permutation on each intersection with a $1$-simplex for which the corresponding homotopy changes $S$ to $P$.
We see that $W_{0}$ is the set of vertices of a normal $2$-sphere $S_{0}=\delta M_{0}$. In $S_{1}$ the pair $u_{1},v_{1}$ will be the ends of a returning arc in at least one $2$-simplex. In another $2$-simplex $u_{1},v_{1}$ will be joined by lines in both $S_{1}$ and $S$ to the vertices of an edge in $S_{0}$. Having such a situation in a $2$-simplex is the only way that removing $u_{1},v_{1}$ will give a normal pattern.
$u_{1}$$v_{1}$xxyywwzz
Figure 9.
Figure 9 shows the situation for $u_{1},v_{1}$ which lie in the $1$-simplex $\gamma_{1}$. In $S_{1}$, $u_{1},v_{1}$ will be the ends of a retuning arc in one of the $2$-simplexes containing $\gamma_{1}$. This cannot be the case in every $2$-simplex containing $\gamma_{1}$ as it would mean that $S_{1}$ would not be connected. In at least one $2$-simplex containing $\gamma_{1}$ as a face, $u_{1},v_{1}$ will be joined to the vertices $w,z$ of an edge of $S_{0}$, shown dashed red. This is the only way that removal of the pair will result in the normal $2$-sphere $S_{0}$. In $P$, $w,z$ will be joined to points $x,y$ by lines, shown blue in Figure 9. Let $\nu_{1}$ be the permutation of $W$, which restricts to $(v_{1},x)(u_{1},y)$ for $W\cap\gamma_{1}$ and is the identity map on all the other $1$-simplexes. Under the corresponding homotopy the two lines become uncrossed if they were crossed before. Note that the points $x,y$ are adjacent and stay adjacent in subsequent homotopies because the homotopy takes place in $M_{0}$ and so the new lines of the homotopy do not intersect the lines of $S_{0}$. The pattern for $S_{0}$ in each $3$-simplex is as in Figure 1 and so the homotopy $\nu_{1}$ must take place in one of the regions which is not bounded by two parallel lines. This is illustrated in Figure 10, which indicates what can happen in the faces of a $3$-simplex. A $4$-track has been replaced by two $3$-tracks. If instead of joining the ordered pair $u_{1},v_{1}$ to the vertices of the second edge, we joined $v_{1},u_{1}$ to those vertices we would get a $6$-strack. To get a subgraph of $P$ one must join up the points to give two $3$-tracks.
If a $2$-simplex $\sigma$ contains $\gamma_{1}$ as one of its faces, but has no intersection with $S_{0}$, then the points will be joined by a returning arc in $\sigma$.
[…]
Figure 10.
We use an induction argument for defining $\nu_{i}$ for $2\leq i\leq k$.
[…]
Figure 11.
Our aim is to show that for each $1$-simplex $\gamma$ we can permute the finite set $\gamma\cap S$ in such a way that under the associated homotopies the spattterned $2$-sphere $S$ becomes the underlying patterned $2$-sphere $P$. It will then be the case that $F$ becomes an isotopy.
Let $\gamma_{i}$ be the $1$-simplex containing $u_{i}$ and $v_{i}$.
We have defined $\nu_{1}$. Let $\mu_{1}=\nu_{1}$. We now define $\nu_{i}$ and $\mu_{i}$ for $i=2,\dots,k$ to be permutations of $W$, where $\nu_{i}$ restricts to a permutation on the points of $W\cap\gamma_{i}$ and is the identity on the other points of $W$, and so that the homotopy associated with $\mu_{i}=\nu_{i}\mu_{i-1}$ moves the points of $W_{i}$ to the positions they should have in $P$. In at least one of the $2$-simplexes containing $\gamma_{i}$ as a face there are lines in $S$ joining $u_{j}$ to $u_{i}$ for some $j<i$ or there is a situation as in Figure 9, in which $u_{i},v_{i}$ are joined to the vertices of an edge of $S_{j}$ for $j<i$. This is because there cannot be a returning arc joining $u_{i}$ and $v_{i}$ in every $2$-simplex containing $\gamma_{i}$, as if there were then $S_{i}$ would not be connected.
If there are lines in $\mu_{i-1}S$ joining $u_{i}$ to $u_{j}$ and $v_{i}$ to $v_{j}$ it will be as in Figure 11(a) or there will be three labelled pairs joined by lines as in Figure 11(b). This follows because $u_{i},v_{i}$ is a removable pair in $S_{i}$ and so it will have a returning arc joining $u_{i}$ and $v_{i}$ and this returning arc was created by one or more homotopy $f_{j}$ for $j<i$.
In Figure 11 (a) we see what happens in a $2$-simplex $\sigma_{i}$ containing $\gamma_{i}$, if the lines joining $u_{i}$ and $v_{i}$ come from a single pair $u_{j},v_{j}$. I think there might be a problem with this Figure. We need both $x,y$ to be in $\gamma_{i}$. It might be the case that both points are in the third edge of the $2$-simplex.
In Figure 11(b) the lines to $u_{i},v_{i}$ come from two pairs. The lines in blue are those of $\mu_{i-1}S\cap\sigma_{i}$. The lines in red are the ones we move them to in $\mu_{i}S$. The bottom edge is the $1$-simplex $\gamma_{i}$ containing the pair $u_{i},v_{i}$ which have labels $i,i$ in the labelling for $S$. In at least one of the $2$-simplexes containing $\gamma$ there will be a $1$-face different from $\gamma$ containing an adjacent pair with labels $l,l$ for $l<i$, which are joined by non-intersecting lines in $P$ to the pair $x,y$ in $\gamma_{i}$, or the situation will be as in Figure 11(b) so that there are pairs with labels less than $i$ in each of the other two faces. The induction step is to apply the permutation $\nu_{i}=(v_{i},x)(u_{i},y)$ to the set $\gamma_{i}\cap W$. Of course if $u_{i}=y$ we apply the transposition $(v_{i},x)$. Again with each transposition we also transpose all attaching lines in all $2$-simplexes containing $\gamma$. This change will not disturb any of the vertices of $W_{i-1}$ or the edges joining them. Note that $x,y$ will have labels bigger than $i$ as if, say, $x\in W_{i-1}$, then in $S_{i-1}$ there would be a point in the intersection of a $1$-simplex with $\mu_{i-1}S$ joined to both $u_{i}$ and $y$ and this does not happen in a spattern.
For each labelled point of $W$ we can choose a line in $\mu S$ joining the point to a point of $W_{0}$ or a point with a smaller label. This will give us a connected subgraph of both $P$ and $\mu S$ that contains every point of $W$. This must mean that $P$ is connected. Since $S$ is a $2$-sphere and $\mu S$ is obtained from $S$ by permuting the points of $W$ and the lines joining them, $\mu S$ is also a $2$-sphere. But we now know $P$ and $\mu S$ have a common spanning subtree. Also $P$ and $\mu S$ have the same number of edges. This means that $P$ is also a $2$-sphere, as $P$ and $\mu S$ will have the same Euler characteristic. The addition of each edge not in a spanning tree results in a region being divided in two, and so there is one extra face for each such edge..
It now follows that at the end of the induction $P=\mu_{k}S$ so that $P$ determines a patterned $2$-sphere.
We now know that $\mu S$ is a patterned surface. All the $2$-pieces, apart from the exceptional one, intersect each $1$-simplex at most once and so are $3$-sided or $4$-sided. The $8$-track shown in Figure 6(a) is the only possibility for the exceptional $2$-piece. Thus $S$ has become an almost normal $2$-sphere and we have a proof of the Poincaré Conjecture. In the homotopy $F$, the first thick sphere $S$ will also be the last thick sphere. After applying $\mu$, then for $j>k$ each step of the homotopy becomes an isotopy in which a returning arc joining adjacent points is removed. There is now a new labelling of $W$ in which every point receives a label $j$, where $n+1\leq j\leq n$. The pair of points labelled $j$ will be a removable pair in $f_{j}(S^{2})$.